5. (Chihara and Hestelberg : Exercise 6.4.25) Let X1, X2, . . . , Xn be random variables with E(X) = μ, for all i-1, 2, . . . , n. Under what condition on the constants ai, a2, . . . , an İs an unbiased estimator of μ?
Answer #1

We have, E(X_i) = \mu , i = 1(1)n.

X = a_1X_1 + a_2 X_2 +....+ a_nX_n

Thus, E(X) = E(a_1X_1 + a_2 X_2 +....+ a_nX_n)

i.e. E(X) = a_1E(X_1)+ a_2 E(X_2) +....+ a_nE(X_n)

i.e. E(X) = a_1\mu + a_2 \mu +....a_n \mu

i.e. E(X) = (a_1+ a_2 +....a_n) \mu

Hence, X will be an unbiased estimator of \mu if,

(a_1+ a_2 +....a_n) = 1.